Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments--Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Molding Fabrication Total
Estimated total machine-hours used 2,500 1,500 4,000
Estimated total fixed manufacturing overhead $ 14,750 $ 17,850 $ 32,600
Estimated variable manufacturing overhead per machine-hour $ 3.30 $ 4.10
Job P Job Q
Direct materials $ 32,000 $ 17,500
Direct labor cost $ 36,200 $ 15,100
Actual machine-hours used:
Molding 3,600 2,700
Fabrication 2,500 2,800
Total 6,100 5,500

weeten Company had no underapplied or overapplied manufacturing overhead costs during the month.

Required:

For questions 1-8, assume that Sweeten Company uses a plantwide predetermined overhead rate with machine-hours as the allocation base. For questions 9-15, assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments.

15. What was Sweeten Company’s cost of goods sold for March?
Answer #1

Molding overhead rate = (14750/2500+3.3) = 9.20 per hour

Fabrication overhead rate = (17850/1500+4.10) = 16 per hour

Calculate cost

Job P Job Q
Direct material 32000 17500
Direct labor 36200 15100
Molding overhead 33120 24840
Fabrication overhead 40000 44800
Total cost 141320 102240

Cost of goods sold = 141320+102240 = $243560